### v1

# M2SVid: End-to-End Inpainting and Refinement for Monocular-to-Stereo Video Conversion

Abstract
We tackle the problem of monocular-to-stereo video conversion and propose a novel architecture for inpainting and refinement of the warped right view obtained by depth-based reprojection of the input left view. We extend the Stable Video Diffusion (SVD) model to utilize the input left video, the warped right video, and the disocclusion masks as conditioning input to generate a high-quality right camera view. In order to effectively exploit information from neighboring frames for inpainting, we modify the attention layers in SVD to compute full attention for discoccluded pixels. Our model is trained to generate the right view video in an end-to-end manner by minimizing image space losses to ensure high-quality generation. Our approach outperforms previous state-of-the-art methods, obtaining an average rank of 1.43 among the 4 compared methods in a user study, while being 6×6$\times$6 × faster than the second placed method.
1 Introduction
Emerging technologies such as Mixed-Reality headsets and glasses allows users to easily experience immersive content, thanks to the use of separate displays for left and right eyes. Rendering a video from left and right eye viewpoints on corresponding left and right displays can create a stereoscopic 3D effect, giving viewers a perceived sense of depth. However, capturing such stereoscopic videos usually requires specialized cameras that can capture both left and right eye perspectives simultaneously. While manual monocular-to-stereo conversion is possible, it is a costly and time-consuming process that requires significant manual effort. Therefore, there is a need for automated, fast and high-quality monocular-to-stereoscopic video conversion methods that can enable large scale conversion of videos.
Figure 1: We present M2SVid, an end-to-end video inpainting and refinement approach for monocular-to-stereo video conversion task. Given an initial right view, obtained via e.g. depth based warping, our method inpaints the missing region and refines the artifacts introduced by warping in a feed-forward manner.
Recent advancements in video generation [5, 4] as well as monocular depth estimation [83, 31] have led to significant improvements in the monocular-to-stereo conversion quality. A common approach is to use depth maps generated by a monocular depth model to warp the input video to a virtual right camera and then inpaint the disoccluded regions [49, 10, 68]. SVG [10] repurposes a pre-trained video diffusion model for the inpainting task, while  [49] trains a custom model. Rather than only modifying discoccluded regions, the recent concurrent work StereoCrafter [68] finetunes a video diffusion model [3] to inpaint and refine the full warped video, using only the warped view and discocclusion masks as input. While obtaining improved results, StereoCrafter still leverages an architecture primarily designed for monocular video inpainting task, without exploiting the constraints specific to stereoscopic video-refinement. Moreover, the inference requires multiple diffusion steps (default 25) leading to high computational cost.
In this work, we propose a novel architecture for efficient inpainting and refinement in the monocular-to-stereo conversion task. Our approach is designed to leverage important additional cues available in this task. Firstly, in the areas that are not disoccluded, the right view largely resembles the left one, where the image content is just horizontally shifted. Thus, even if depth-based warping introduces artifacts in the generated right view, we conjecture that most of them can be fixed by relying on left view information. Secondly, the disoccluded regions are generally small in most cases. Furthermore in the presence of camera or scene motion, the disoccluded regions in one frame are often visible at nearby frames, thus simplifying the inpainting task into a spatial-temporal matching problem. The aforementioned aspects motivate us to design M2SVid, a specialized architecture which performs inpainting and refinement in an feed-forward manner, without any iterative diffusion steps.
Our main contributions are the following. First, we extend the Stable Video Diffusion (SVD) architecture by employing the input left view, in addition to the warped right view and disocclusion mask as conditioning for the refinement task. This makes it easier for the model to propagate information from the uncorrupted left view to the target right view, thus preserving high frequency details such as fine structures. Secondly, rather than performing temporal attention only over the same spatial locations as in SVD [3], we compute full cross-attention for disoccluded pixels. This provides greater flexibility to propagate useful information from either nearby temporal frames or the associated left view, with a limited computational overhead. Finally we train our model in an end-to-end manner using perceptual and fidelity based losses, enabling high quality inpainting and refinement (See Fig. 1).
We perform quantitative evaluation, qualitative analysis, as well as a human perception user study to evaluate our method. It shows that our M2SVid outperforms previous state-of-the-art StereoCrafter and SVG, obtaining an average rank of 1.43 / 4.0 in the user study, while being 6×6$\times$6 × and 600×600$\times$600 × faster than StereoCrafter and SVG respectively.
2 Related Work
2.1 Novel View Synthesis
Early methods for novel-view synthetis performed 3D reconstruction of the full scene given dense input views [65, 28, 84], which can then be used to render novel views. While Multi-Plane Image (MPI)-based approaches [74, 18, 73] were successfully used to generate multiplane representations from single images, NeRFs [52] marked a milestone in the field, and spawned many efforts to improve quality [1, 2], reduce latency [54, 19, 78, 85, 55], extend to large-scale scenes [75], explicit representation [32], pose-free setup [40, 77, 8, 51, 81, 72], or fewer inputs [23, 33, 56, 11, 59, 72, 80], including just a single image [50, 42, 43, 44]. Approaches such as CAT3D [15], NVS-Diffusion [29], MultiDiff [53] directly generate the queried camera view using a diffusion models. A few works also extend these to videos [79]. While obtaining promising results, the generated views can still have artifacts due to the inherent difficulties of rendering views with large viewpoint changes.
2.2 Monocular-to-Stereo Conversion
Compared to the general novel-view synthesis problem, Monocular-to-Stereo conversion aims to only render a fixed right camera view given an input left view. This can allow viewing any monocular content in 3D on mixed reality headsets, enabling a more immersive experience. The majority of previous works can be categorized into two families. One line of work aims to directly generate the right image/video given the left input [82, 82, 67, 36]. A notable example is Deep3D [82], which leverages an internal disparity representation to directly render the right video given the left one. The second type of approaches employ an explicit depth map to reproject the left image to the right camera, and then inpaint the disoccluded regions [69, 24, 49, 10, 68, 35, 76, 10]. Notably, Wang et al. [76] learn a diffusion model for inpainting disoccluded pixels using self-supervision with random cycle rendering. while SVG [10] uses a pre-trained video diffusion model for inpainting. While the method can achieve spatio-temporal consistency in the output, the inpainting result can be incorrect due to the lack of task-specific fine tuning. Mehl et al. [49] aim to instead perform a geometry aware inpainting of the disoccluded areas using local background information, rather than generating arbitrarily realistic content. StereoCrafter [68], which is concurrent to our work, fine-tunes a video diffusion model for inpainting the disoccluded areas, on a large stereo dataset (not released). Another concurrent work SpatialDreamer [47] mitigates the necessity of paired stereoscopic training data by proposing a self-supervised training framework using a forward-backward rendering mechanism. Note that a number of the above methods do not release code or model, and are often trained on private collected data, making a fair comparison difficult.
2.3 Diffusion Models
Diffusion Models [20, 70, 60] are a family of generative models that iteratively denoise an image to generate the output. They have been tremendously successful for the text-to-image [20, 60, 63, 87, 58] and text-to-video [4, 21] tasks. They have also been successfully employed for a wide variety of computer vision tasks, thanks to the strong image priors learned by the models. These include image generation applications such as image [46, 9, 41] and video [92, 7, 37] inpainting, novel view synthesis [53, 27], as well as other pixel-wise prediction tasks such as monocular depth estimation [31, 22, 16, 66], semantic segmentation [30], or normal estimation [14]. Conventionally, diffusion models employ multiple denoising steps during inference, leading to large computations cost. A number of approaches aim to train one-step models to address this, using knowledge distillation [86, 90], GAN training [48], low-rank adaptors [90]. Garcia et al. [16] instead finetune a pre-trained diffusion model to perform direct feed-forward monocular depth estimation using end-to-end training. In this work, we show that such an end-to-end training strategy can also be employed for inpainting disoccluded regions in monocular-to-stereo video conversion task.
3 Conditional Latent Diffusion Models
Here, we provide a brief background on diffusion models employed in our method. Denoising Diffusion Probabilistic Models (DDPMs) [20] are generative models trained to map a simple noise distribution $p_{T}$ to the data distribution $p_{0}$, by reversing a stochastic forward process $p_{t}$, $t=1,...,T$. A denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t)$ is trained to generate an image from pure noise by progressively denoising inputs $x_{t}$ at time stamp $t$. In order to reduce computational complexity, Latent Diffusion Models (LDMs) [60] operate in a latent space of a Variational Autoencoder (VAE) [34]. Conditional diffusion models [62, 88] extend the denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t,c)$ to be conditioned on additional input $c$, such as text [62], images [62], etc. to control the generation process.
Training and inference with conditional LDMs:  During training, given a data sample $\mathbf{x}$ (e.g., an image or video) and its corresponding conditioning input $c$ (e.g., text or another image), the data sample $\mathbf{x}$ is first encoded into a latent representation $\mathbf{z}=E(\mathbf{x})$ using VAE encoder $E$. The latent representation $\mathbf{z}$ is perturbed through a forward diffusion process: $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ is Gaussian noise, and $\bar{\alpha}_{t}$ controls the noise schedule. Instead of predicting the noise $\boldsymbol{\epsilon}$ directly, diffusion models usually leverage v-parameterization [64], where the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{z}_{t},t,c)$ learns to predict: $\boldsymbol{\mathbf{v}}=\alpha_{t}\boldsymbol{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}.$ Therefore the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$, typically a U-Net [61], is trained to reconstruct $\boldsymbol{\mathbf{v}}$, by optimizing the objective:
$\mathcal{L}=\mathbb{E}_{\mathbf{z},c,\boldsymbol{\epsilon}\sim\mathcal{N}(0,I),t\sim\mathcal{U}(T)}\left[\|\boldsymbol{\mathbf{v}}-\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{z}_{t},t,c)\|^{2}\right].$ (1)
During inference, the denoising process starts from pure noise $\boldsymbol{\mathbf{z}}_{T}$ and the learned denoiser $\boldsymbol{\mathbf{v}}_{\theta}(z_{t},t,c)$ iteratively refines the output over $T$ steps to recover the final sample.
Diffusion as feed-forward models:  Garcia et al. [16] propose to use pre-trained diffusion U-Net model as a feed-forward models for pixel-to-pixel tasks such as depth and normal estimation. In this scenario, the timestep $t$ is not sampled anymore and fixed as $t=T$ to train the model for single-step prediction. The noise $\boldsymbol{\epsilon}$ is additionally replaced by the mean of the noise distribution, i.e., zero. With $t=T$, we get $\bar{\alpha}_{T}\approx 0$, and thus $\mathbf{z}_{T}=\sqrt{\bar{\alpha}_{T}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{T}}\boldsymbol{\epsilon}\approx\mathbf{0}$ and $\boldsymbol{\mathbf{v}}=\alpha_{t}\boldsymbol{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}\approx-\mathbf{z}.$ As such, the model directly learns to predict the clean output $\boldsymbol{\mathbf{v}}\approx-\mathbf{z}$ starting from zero noise vector $\mathbf{z}_{T}=\mathbf{0}$. With such timestep sampling and assumptions, the standard loss (Eq. 1) converges to the $L_{2}$ loss in the latent space between the ground truth $-\mathbf{z}$ and the predicted latents $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{0},T,c)$:
$\mathcal{L}_{latent}=\mathbb{E}_{\mathbf{z},c}\left[\|((-\mathbf{z})-\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{0},T,c)\|^{2}\right].$ (2)
However, [16] proposes decoding the predicted latents, $\mathbf{\hat{z}}=-\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\mathbf{0},T,c)$, using a VAE decoder $D$ to obtain the reconstructed output $\mathbf{\hat{x}}=D(\mathbf{\hat{z}})$. Instead of minimizing the latent space error Eq. 2, they propose to optimize a task-specific loss directly in image space with $\mathcal{L}=L_{task}(\mathbf{x},\mathbf{\hat{x}})$.
Figure 2: Overview of our monocular-to-stereo conversion pipeline. Given an input monocular video, we first estimate per-pixel depth, which is used to warp the input video to a right camera view. The input video, the warped video, as well as the disocclusion masks are then passed to our video inpainting and refinement module to generate the final right view.
4 Monocular-to-Stereo Conversion Pipeline
In this work, we tackle the problem of converting any 2D video to a 3D stereoscopic video. Given a monocular video $V\in\mathbb{R}^{N\times H\times W\times 3}$, containing N frames of resolution $H\times W$, the goal is to transform it into a stereoscopic video pair $(V_{l},V_{r})$, where $V_{l},V_{r}\in\mathbb{R}^{N\times H\times W\times 3}$ represent the views corresponding to the left and right eyes, respectively. Following prior work, we assume that the input video $V$ corresponds to the left-eye view, i.e., $V_{l}=V$, and therefore, the goal is to only generate the right-eye view $V_{r}$ from $V$. Note that multiple possible right-eye views $V_{r}$ exist depending on the baseline, i.e. the distance between cameras. Hence we condition the generation on a maximum desired disparity $D_{\text{max}}$ between $V_{l}$ and $V_{r}$. This can be provided by the user to control the amount of stereoscopic effect in practical applications.
An overview of the pipeline is shown in Fig. 2. We follow the commonly employed strategy [49, 68] of performing depth-based reprojection to generate an estimate of the right view, as described next.
Video-depth estimation:  We use a monocular video depth estimation model to obtain per-frame depths. Our method is agnostic to the choice of depth estimation method.
Left-to-right video warping:  We use the user-provided maximum disparity $D_{\text{max}}$ to convert the depth maps to disparities. The disparity maps are then used to forward-splat the input left-view video to a virtual right view denoted as $V_{r}^{warp}$. Note that the warped right frames inevitably contain missing values (i.e. holes) introduced by disocclusions. These missing pixels are indicated by the disocclusion mask $M_{r}^{occ}$, obtained as a by-product of the splatting.
Video inpainting and refinement:  The final step in our pipeline is to generate a high quality right view given the warped video $V_{r}^{warp}$. As mentioned before, the warped right video $V_{r}^{warp}$ contains holes introduced by disocclusions. Furthermore, the warped right video $V_{r}^{warp}$ can also contain artifacts introduced due to reprojection and depth errors. We use a Video Diffusion Model to inpaint the missing regions, as well as correct the artifacts present in the warped right views.
In this work, we mainly focus on the final “video inpainting and refinement” task. To this end, we introduce an efficient end-to-end model to inpaint the disoccluded regions and fix reprojection artifacts. Our architecture is described in detail in the next section.
5 End-to-End Stereoscopic Video Refinement
We introduce M2SVid, an efficient end-to-end model for generating a high-quality right video from the reprojected right video. While this problem is similar to the standard video inpainting task, it comes with a set of unique properties, as described next.
(1)  In our setup where we render a right camera view, the inpainting holes are always introduced in regions where there is an increase in the depth value, as we move horizontally from left to right in the image. Furthermore, the inpainting regions are guaranteed to correspond to the background, i.e. the far plane. Thus, as opposed to a general-purpose inpainting model such as for object removal, our inpainting model can learn to use context from only one side of the inpainting regions, corresponding to the far plane. (2)  In a general inpainting setup, the model is only required to generate the missing pixels, while keeping the rest of the image unaltered. However in our case, the regions outside the inpainting mask can also contain artifacts introduced by errors in the depth map, or interpolation during forward splatting. These artifacts are especially common in presence of thin structures such as fences, where the estimated depth can have substantial errors. Fortunately, we can leverage the original left video to correct the errors introduced by the reprojection operation.
(3)  In a single image inpainting task it is necessary to hallucinate content in the missing regions due to lack of additional information. A video inpainting model can leverage information from multiple frames. However when the inpainting holes are large, the model is still required to perform substantial hallucination. In contrast, the inpainting regions are expected to be much thinner in our task (i.e., as controlled by the maximum disparity $D_{\text{max}}$ parameter). As a result, the inpainting problem is greatly simplified and the model can try to copy information from other temporal frames to avoid hallucination.
We develop a custom architecture for the right view generation task aiming to exploit the aforementioned properties.
Figure 3: An overview of our proposed stereoscopic video refinement method. Our model inpaints the discoccluded regions in the warped right view, and corrects possible artifacts introduced by warping errors. The model takes the VAE encodings of the input left view, reprojected right view, and the disocclusion mask as conditioning to the U-Net. The latent encodings of the refined right view are then generated in a single denoising step, and then decoded by the VAE Decoder to generate the output right video. In order to effectively utilize the information from neighboring frames for inpainting, we extend the spatial attention layer in SVD to compute full attention for the disoccluded tokens. The model is training end-to-end by minimizing image space and latent space losses.
5.1 Overview
An overview of our method is shown in Fig. 3. We base our model on a strong pre-trained video diffusion model to benefit from learned video priors. In particular, we utilize Stable Video Diffusion (SVD) [3], a latent video diffusion model trained to generate videos from an input image. We customize the SVD architecture for the stereoscopic video generation task as follows. First, we condition each latent generation on the input left video $V_{l}$, reprojected right video $V_{r}^{warp}$, as well as the disocclusion masks $M_{r}^{occ}$. Secondly, only for the inpainted pixels, we extend the spatial attention in SVD to operate over all pixels in all frames, instead of just the pixels in the same spatial location. This gives greater flexibility for the model to copy visible pixels from neighboring frames in order to consistently inpaint. Our model is trained end-to-end on stereoscopic datasets to generate the refined right view in a feed-forward manner.
5.2 High-Frequency Details Preservation
Conventional video inpainting methods utilize masked video to condition inpainting models, a strategy widely adopted in stereo conversion methods [68, 49, 76], where the inpainting model is conditioned solely on the warped right video $V_{r}^{warp}$ and the disocclusion mask $M_{\text{r}}^{\text{occ}}$. However, as discussed in Section 5 (1), the warped right video is likely to contain artifacts even in non-inpainting regions. Thus we propose to also utilize the original left video $V_{\text{l}}$ as an extra input to the model. In more detail, the standard SVD model takes the noisy latents for a video snippet and the VAE encoding of a conditioning image as input. We modify the SVD architecture to instead take the following inputs as conditioning: the VAE encoding of the left video snippet $E(V_{l})$, the VAE encoding of the warped right snippet $E(V_{r}^{warp})$, and the disocclusion mask $M_{\text{r}}^{occ,resized}$, resized to the same resolution as the VAE encodings. We thus denote the model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\boldsymbol{\mathbf{z}}_{t},t,c)$ with conditioning inputs $c$ as
$c=[E(V_{l}),E(V_{r}^{warp}),M_{\text{r}}^{occ,resized}]\,\,,$ (3)
where [.][.][ . ] refers to concatenation. This modification is achieved by modifying the first convolution layer in the U-Net, as is the common practise.
Using the original left video as conditioning allows to model to easily infer high frequency details as well as other information which might have been destroyed during the depth based reprojection stage. As shown in Fig. 6, this improves the quality of the generated right views, with high-frequency details from the left view better preserved.
5.3 Spatio-Temporal Aggregation for Inpainting
The SVD model is trained to take multiple video frames (i.e. a snippet) as input and jointly denoise them in order to generate a temporally consistent output. Ideally, one would like to compute a full attention over all spatial tokens in all frames to enable maximal information flow. However, this operation would be prohibitively expensive, resulting in a complexity of $H=N^{2}\times h^{2}\times w^{2}$ for each attention layer, where $h$ and $w$ are the size of the latent representations. For this reason, the standard video diffusion models such as SVD and VideoCrafter [5] only factorize the full attention computation into interleaved spatial and temporal attention layers, to propagate information across multiple locations and frames. In the spatial attention layers, only tokens from the same timestamp attend to each other, while in the temporal attention layers, tokens from the same spatial location attend to each other across different timestamps. This drastically reduces the complexity to the order of only $N\times h^{2}\times w^{2}+N^{2}\times h\times w$, which is more palatable.
The factorization of the full attention computation is however likely to reduce the modelling capacity for our task, specially in dynamic scenes with camera motion. In such cases, as discussed in Sec. 5 (2), an occluded pixel in one frame is likely visible at a different spatial location in another frame, simplifying the inpainting problem. In such cases, we would like to perform a full spatio-temporal attention. Fortunately, the number of inpainted tokens constitute only a small fraction of all tokens (usually $<$ 5%). We exploit this property and modify the spatial attention layer in SVD to allow tokens corresponding to the disocclusion mask $M^{occ}_{r}$ to attend to all other tokens, while using the spatial attention only for other tokens. This improves the inpainting capability of our model (Fig. 5) without any significant computational overhead.
5.4 Efficient Feed-Forward Prediction
The diffusion models for text-to-image/video generation and inpainting commonly utilize multiple denoising steps to generate the final image [20, 62, 60]. This is important due to the ill-posed nature of the problem wherein multiple solutions exists for a single input prompt. Utilizing fewer denoising steps usually results in blurry output.
However, in our task of right view generation, the inpainting regions are generally small. Furthermore, as described in Sec. 5, the neighboring temporal frames as well as the input left frames contain a significant amount of information needed to generate a high-quality, temporally consistent right video. The model therefore does not need to perform significant hallucination in most cases, but rather needs to fetch the relevant information from other frames. The relatively constrained nature of our problem thus motivates us to generate the right view in a direct feed-forward manner, similar to the approach employed by Garcia et al. [16] for depth estimation.
Given the input left video $V_{\text{l}}$, warped right video $V_{\text{r}}^{warp}$, and the disocclusion masks $M_{\text{r}}^{occ}$, we generate the conditioning input $c$ to the model by obtaining the VAE encodings of left and reprojected videos, as shown in Eq. 3. The initial latent is set to the mean noise 0 as in [16]. The conditioning input, together with the initial latent are then passed through the Video U-Net to obtain the latent encoding $\hat{z}$ of the right video. This is then passed through the VAE decoder $D$ to obtain the predicted right view $\hat{V}_{r}$. The inference pipeline can thus be denoted as,
$\hat{V}_{r}=D(-\hat{\boldsymbol{\mathbf{v}}}_{\theta}(\textbf{0},T,c))$ (4)
Our feed-forward prediction strategy significantly reduces the latency of the right view generation step compared to existing methods, as shown in Table 2. Furthermore, this also allows us to train the model end-to-end using image quality losses w.r.t. to the ground truth right video $V_{r}$, in order to maximize the quality of the generated video. In particular, we train the model using a combination of LPIPS and $L_{\text{1}}$ losses directly in the image space, along with an auxiliary loss Eq. 2 in the latent space. Our final training loss is thus defined as,
$\mathcal{L}=\mathcal{L}_{latent}(z,\hat{z})+\mathcal{L}_{\text{L1}}(V_{r},\hat{V}_{r})+\mathcal{L}_{\text{LPIPS}}(V_{r},\hat{V}_{r})$ (5)
Here $z$ corresponds to the VAE encodings of the ground truth right video $V_{r}$. Note that during the training, keep the VAE decoder frozen and only fine tune the U-Net model.
5.5 Training Data
In order to train our models using the supervised objective Eq. 5, we require datasets consisting of stereoscopic videos, together with the disparity maps between the stereo pairs. Due to the lack of large-scale stereoscopic datasets, prior works have focused on collecting and processing private datasets from the internet, which usually have not been publicly released. Instead we aim to train using publicly available datasets.
We utilize Ego4D [17] and Stereo4D [25] datasets for our training. Ego4D a large-scale egocentric video dataset containing approximately 263 long videos (80 hours in total) collected using a stereo camera, while the recently released Stereo4D [25] dataset consists of $\sim$ 200K stereoscopic video clips sourced from $\sim$ 7K online videos. The Stereo4D dataset also provides rectified videos and disparity maps. The Ego4D dataset on the other hand contains unrectified videos, without any precomputed disparity maps.
We thus perform the following steps to preprocess Ego4D dataset in order to use it for training. First, we uniformly sample $K=200$ frames from the video and perform dense feature matching using LoFTR [71]. These are used to compute the fundamental matrix [91] using RANSAC [12]. Next, we estimate the rectification transformations for both left and right frames. After rectifying the images, we spatially crop the video to discard boundary regions with missing pixels. We then compute the LoFTR matches again and shift the left and right views horizontally until all disparities between the matched points are positive and the smallest disparity is zero. This ensures that the left and right videos follow a rectified stereo setup. Finally, we utilize an off-the-shelf stereo matching method BiDaVideo [26] to obtain dense disparity maps for all stereo pairs.
Figure 4: Qualitative comparison of our approach with state-of-the-art methods SVG [10] and StereoCrafter [68]. Our approach can effectively preserve the high-frequency information from the input video and generate high-quality right views.
6 Experimental Results
In this section, we evaluate the quality of the right view videos generated by our method, both qualitatively as well as quantitatively. Further results, analysis, visualizations and implementation details are provided in the Appendix.
Quantitative evaluation:  To quantitatively evaluate our approach, we need datasets containing left and right stereo videos, together with the (pseudo) ground-truth disparity maps which are used to generate the warped right views for each method. This allows us to directly compare the generated right views with the ground truth right views, without having to worry about errors introduced by incorrect depth estimation. We rely on the test sets of the Stereo4D [25] and Ego4D [17] datasets, processed as described in Sec 5.5. As metric, we use the standard image quality metrics PSNR, MS-SSIM, and LPIPS [89] to compare the different methods. We compute the metrics independently over each video, and average them over the full dataset. In the appendix, we also report the metrics computed only inside, as well as outside the disoccluded regions. All evaluations are performed on 16-frame videos, sampled at 5 FPS, resized to 512 resolution, and centrally cropped.
Qualitative evaluation:  For qualitative analysis and user studies, we only need monocular videos along with per-frame depth maps. We use videos from the DAVIS dataset [57] as well as videos from free online stock sources [^1] https://www.pexels.com/.We sample 16 frames per video at 8 FPS and resize them to $640\times 1152$ resolution. We compute the depth maps using the recent DepthCrafter [22] method.
Method Training data Denoising steps PSNR $\uparrow$ MS-SSIM $\uparrow$ LPIPS $\downarrow$
SVG [10] - (training free) 50 steps 25.6 0.926 0.217
StereoCrafter [68] private dataset 25 steps 24.9 0.909 0.242
StereoCrafter [68] private dataset 1 step 25.3 0.911 0.262
M2SVid (Ours) Stereo4D + Ego4D 1 step 26.2 0.915 0.180
Table 1: State-of-the-art comparison on Stereo4D test set. Our approach obtains the best scores in terms of PSNR as well as LPIPS.
6.1 Comparison to State-of-the-Art
Baselines:  We compare our method to the recent state-of-the-art monocular-to-stereo conversion models SVG [10] and StereoCrafter [68]. SVG is a training-free method that utilizes a frozen diffusion model to inpaint regions within the disocclusion mask while not modifying the remaining regions. StereoCrafter, which is an unpublished concurrent work, fine-tunes SVD [3] to inpaint and refine the warped right views in a diffusion-based denoising setup, using the warped right views along with the disocclusion masks as conditioning. We evaluate StereoCrafter using the default 25 denoising steps as well as single denoising step. We exclude the Deep3D [82] from the evaluation as it doesn’t allow control over the baseline between the stereo cameras and is shown to have inferior results [68, 10] compared to SVG and StereoCrafter. In order to ensure a fair comparison, we use the same depth maps for reprojection in all methods. However, we use the official warping implementation provided by the authors for each method.
Qualitative Results:  We perform a qualitative comparison with the SVG and StereoCrafter methods in 4. Since SVG only performs inpainting within the disocclusion mask, it fails to fix the artifacts introduced by errors in warping. Furthermore, it can incorrectly extend the ‘foreground’ object to fill the inpainting hole due to lack of task-specific fine tuning. StereoCrafter with a single denoising step produces blurry inpainting and degrades the quality of the warped areas as well. This is expected since the model was trained using a multi-step denoising objective. When using 25 denoising steps, StereoCrafter inference generally produces sharp results. However it can struggle at times to correctly generate the high-frequency details. This is because the high-frequency information can often get degraded during the warping step. Since StereoCrafter only utilizes warped view and masks as conditioning, it can struggle to recover the details. In contrast, our model M2SVid is conditioned on the input left view as well, allowing it to leverage the full context for inpainting and refinement. Furthermore, since our method is trained in an end-to-end manner with image space losses, it can learn to minimize the loss of high-frequency information introduced by VAE decoder.
Method Denoising steps Average rank $\downarrow$ Latency (s) $\downarrow$
SVG [10] 50 steps 2.88 1270.4
StereoCrafter [68] 25 steps 2.05 12.2
StereoCrafter [68] 1 step 3.46 2.4
M2SVid (Ours) 1 step 1.43 2.1
Table 2: Average human perception study rankings over 21 videos with 13 distinct users. The ranking ranges from 1 to 4, with 1 being the best. Our method significantly outperforms StereoCrafter with an average rank of 1.43. We also show the run-times computed using an A100 GPU on a $512\times 512$ videos with 16 frames.
Quantitative Results:  We quantitatively compare our approach with SVG and StereoCrafter on Stereo4D in Tab. 1. Despite our method M2SVid performing only a single step inference, it significantly outperforms StereoCrafter on all metrics and obtains the best PSNR and LPIPS scores.
User study:  We also perform a user study to evaluate the quality of the generated right views. We use 21 videos containing 16 frames each for the study. The videos are sourced from DAVIS dataset, as well as publicly available sources. For each video, we show the right videos generated by each method (anonymized). Additionally, we also show the input left video, and the disocclusion mask as reference. Each of the 13 participants were asked to rank the quality of the generated videos from 1 (best) to 4 (worst), taking into account factors such as temporal consistency, image quality (i.e. sharpness) and lack of artifacts. We collected 112 rankings per method. The average rankings are summarized in Tab. 2. Our method significantly outperforms all others, obtaining an average ranking of 1.43, compared to 2.05 obtained by StereoCrafter with 25 denoising steps.
Figure 5: Using full-attention at dis-occluded pixels (Sec. 5.3) enables the model to exploit information from visible pixels in other frames to improve inpainting.
Figure 6: Impact of our left view conditioning (Sec. 5.2).
Run-time:  According to Tab. 2, on an A100 GPU, our approach M2SVid achieves a 6$\times$ and 635$\times$ speed-up compared to StereoCrafter with 25 steps and SVG respectively.
6.2 Ablation Study
We ablate the key components of our approach in this section on the Ego4D and Stereo4D datasets.
Left view conditioning (Sec. 5.2):  The impact of using the left view as conditioning, in addition to using the warped view and disocclusion mask is shown in Tab. 3. Conditioning on the left video shows a 6.0%, 5.8% and 9.8% improvement in PSNR, MS-SSIM and LPIPS metrics respectively on Ego4D with 25 denoising steps. A similar improvement is observed in the feed-forward case. This is because the left video conditioning allows the model to recover high-frequency details and correct artifacts introduced during the warping, as seen in Fig. 6.
Model feed Inf. Stereo4D Ego4D
[…]
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ 25 24.5 0.886 0.226 20.8 0.822 0.296
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ + $V_{l}$ 25 24.8 0.891 0.215 22.1 0.870 0.267
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ 1 26.1 0.913 0.187 21.5 0.837 0.276
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ + $V_{l}$ 1 26.2 0.915 0.179 22.8 0.886 0.244
+full attention 1 26.2 0.915 0.180 22.7 0.885 0.248
Table 3: Impact of left-view conditioning and full attention at disoccluded pixels on Stereo4D and Ego4D datasets.
Loss feed Inf. Stereo4D Ego4D
[…]
A Standard loss (1) 25 24.8 0.891 0.215 22.1 0.870 0.267
B Standard loss (1) 1 25.7 0.901 0.242 23.0 0.877 0.320
C $L_{latent}$ (2) 1 25.6 0.901 0.238 22.9 0.875 0.318
D $L_{latent}+L_{\text{LPIPS}}+L_{\text{L1}}$ (5) 1 26.2 0.915 0.179 22.8 0.886 0.244
Table 4: Impact of different inference strategies and losses.
Full-attention at disoccluded pixels (Sec. 5.3):  While this contribution doesn’t impact metrics in Tab. 3, in Fig. 5, we can observe that computing full-attention in the inpainting regions prevents hallucination of thin structures.
End-to-end training (Sec. 5.4):  In Tab. 4, we ablate our training strategy. When comparing the model trained with the standard diffusion loss (Eq. 1) with 25 (A) or 1 (B) diffusion step at inference, the single step model leads to more blurry results than (A), as evidenced by the 12% and 20% improvement in LPIPS on Stereo4D and Ego4D respectively. End-to-end training with only latent space supervision (Eq. 2) (C) only slightly improves the LPIPS metric. Our final training loss (Eq. 5) (D), including LPIPS and L1 losses in image space, largely improves the image sharpness, as evidenced by the 24% and 22% improvement in LPIPS on Stereo4D and Ego4D compared to (C). Notably, we outperform the standard diffusion loss with 25 inference steps (A) on all metrics and datasets.
7 Conclusion
We introduce an end-to-end approach for stereoscopic video inpainting and refinement in this work. We utilize Stable Video Diffusion as our base model, and extend it to input the warped right view, discocclusion mask, and the input left view as conditioning. Next, we modify the attention layers in SVD to compute full attention for the discoccluded pixels in order to improve inpainting quality. Crucially, we perform the video refinement using a single denoising step, enabling end-to-end training with image space losses. Qualitative as well as quantitative experiments demonstrate that our method clearly outperforms prior state-of-the-art methods for the monocular-to-stereo video conversion task.
Supplementary Material
The supplementary material provides more details about the datasets, implementation, and results. Appendix A provides a brief background on Diffusion Models. More details about our training and evaluation datasets are provided in Appendix B. Appendix C provides further implementation details, while Appendix D contains more detailed results and analysis. Finally, Appendix E provides details on inference for high-resolution long videos. We also include the results generated by our methods as well as the compared approaches, on the 21 videos used for our user study. You can view each of them independently, or open the index.html file in a browser. We also provide additional results generated by our model on high-resolution long videos as described in Appendix E.
Appendix A Background on Diffusion Models
DDPMs:  Denoising Diffusion Probabilistic Models (DDPMs) [20] are generative models trained to map a simple noise distribution $p_{T}$ to the data distribution $p_{0}$, by reversing a stochastic forward process $p_{t}$, $t=1,...,T$. This forward process gradually adds small amounts of Gaussian noise with variance $\beta_{t}$, ensuring the reverse process can be approximated as a Gaussian distribution. The forward process might be also defined by $\boldsymbol{\mathbf{x}}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\mathbf{\epsilon}}$, where $\boldsymbol{\mathbf{x}}_{0}$ is a data sample, $\boldsymbol{\mathbf{\epsilon}}\sim\mathcal{N}(\text{0},\text{I})$ is a noise, and $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}\alpha_{\tau}$. A denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t)$ is trained to progressively remove noise from the input, by predicting $\boldsymbol{\mathbf{x}}_{t-1}$ from $\boldsymbol{\mathbf{x}}_{t}$ and timestamp $t$, effectively reversing the diffusion.
DDIMs:  Denoising Diffusion Implicit Models (DDIMs) [70] extend DDPMs by introducing non-Markovian diffusion processes, preserving the original training objective while enabling much faster inference with as few as 25 or 50 steps.
LDMs:  To facilitate training, Latent Diffusion Models (LDMs) [60] operate in the latent space of a Variational Autoencoder (VAE) [34], which consists of an encoder $E$ and a decoder $D$. The VAE is trained separately and remains frozen during diffusion model training. By working in the latent space, where $D(E(x))\approx x$, LDMs significantly reduce computational complexity while preserving essential data structure. The diffusion model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}$ is then trained directly in this lower-dimensional latent space.
Conditional diffusion models:  Finally, conditional diffusion models [62, 88] extend diffusion models by conditioning the denoising model $\hat{\boldsymbol{\mathbf{v}}}_{\theta}(x_{t},t,c)$ on additional input $c$, such as text [62], images [62], or pose and depth maps [88].
Appendix B More Details on Datasets
B.1 Stereo4D
Stereo4D: The dataset contains $\sim 200$K video clips sourced from $\sim 7.5$K online videos. Each clip lasts between 2–6 seconds with a spatial resolution of $512\times 512$. The videos cover diverse scenes featuring both indoor and outdoor activities. While the dataset is relatively large, it also includes many static videos or videos with low disparity. We hold out a random subset of 400 videos for method evaluation and use the rest for the training. We source only one clip per video that is longer than 3.2 seconds and sample 16 frames at 5 fps, resulting in a test set of 400 diverse video clips with a resolution of $16\times 512\times 512$.
B.2 Ego4D
Ego4D: This is one of the largest egocentric video datasets, where participants were asked to perform various activities. While most videos are monocular, the dataset includes 263 long stereoscopic videos (80 hours in total) with a resolution of $1400\times 1400$. As the videos are egocentric, captured using a head-mounted camera, they exhibit significant camera movement and contain close objects with a large disparity range between the stereo views. We split 263 stereo videos of the dataset into $\sim$ 57k 5-second video clips. We hold out all clips from one video, namely 200 clips, for method evaluation, while using the rest 57K clips for training.
Ego4d Pre-processing:  We perform rectification for each long video as described in Sec. 5.5 of the main paper. For video rectification, we uniformly sample 200 frames from each video. Then, we split the videos into 5-second subsequent video clips and compute shifts and disparity at the clip level. Finally, to ensure high-quality rectified data, we perform LoFTR feature matching [71] again and filter point pairs by computing the fundamental matrix with RANSAC [13], applied to frames sampled at 1 FPS from each clip. We then filter out clips where the matched points exhibit more than 2 pixels of vertical disparity.
B.3 DAVIS
DAVIS: The dataset consists of 50 videos, each up to 6 seconds long, with a resolution of $1024\times 1920$. The videos are highly dynamic, featuring moving cameras and one or more moving objects (e.g., BMX riding, breakdancing, etc.). Note that the dataset only contains monocular videos, without an ground truth depth or right view. Hence we use this dataset solely for qualitative comparisons and user studies.
Figure 7: The interface shown to the users during the human perception study. On the left, we showed the input left video, together with the disocclusion mask. For each video, we showed results of the four method. Note that the users were not shown any labels to indicate which outputs were generated by which method. Furthermore, we randomized the order in which the methods were listed for each video. The users were asked to assign rankings from 1 (best) to 4 (worst) for each method. In cases two or more methods were indistinguishable, the users were given the choice of assigning the same rank to multiple methods. The users could play or pause the videos as they wish, view the video in full screen, or change the playback speed if they desired.
Figure 8: The users were shown the above intructions during the study.
Method Training data Denoising steps Full view Inside disocclusion mask only Outside disocclusion mask
[…]
SVG [10] - (training free) 50 steps 25.6 0.926 0.217 38.1 0.994 0.014 26.3 0.940 0.190
StereoCrafter [68] private dataset 25 steps 24.9 0.909 0.242 38.2 0.995 0.012 25.5 0.922 0.217
StereoCrafter [68] private dataset 1 step 25.3 0.911 0.262 38.9 0.996 0.014 25.8 0.923 0.234
M2SVid (Ours) Stereo4D + Ego4D 1 step 26.2 0.915 0.180 38.2 0.994 0.010 26.8 0.924 0.161
Table 5: Quantitative comparison of our approach with state-of-the-art methods on Stereo4D test set, in terms of PSNR, MS-SSIM, and LPIPS. We report the metrics computed over the full image, only inside the discoccluion mask, and only outside the disocclusion mask.
Model End-to-end Denoising Full image Inside disoccluded regions
Architecture training steps Stereo4D Ego4D Stereo4D Ego4D
[…]
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ 25 steps 24.5 0.886 0.226 20.8 0.822 0.296 36.1 0.993 0.011 29.3 0.986 0.027
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ + $V_{l}$ 25 steps 24.8 0.891 0.215 22.1 0.870 0.267 36.7 0.994 0.011 30.1 0.988 0.024
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ 1 step 26.1 0.913 0.187 21.5 0.837 0.276 38.0 0.994 0.010 30.4 0.987 0.026
Cond. with $V^{warp}_{r}$ + $M^{occ}_{r}$ + $V_{l}$ 1 step 26.2 0.915 0.179 22.8 0.886 0.244 38.2 0.995 0.009 30.9 0.989 0.024
+full attention 1 step 26.2 0.915 0.180 22.7 0.885 0.248 38.2 0.994 0.010 30.7 0.989 0.025
Table 6: We analyse the impact of different conditioning inputs as well as inference modes on Stereo4D and Ego4D datasets. The metrics are reported over the full images, as well as only the disoccluded regions.
Loss End-to-end Inference Full image Inside disoccluded regions
training steps Stereo4D Ego4D Stereo4D Ego4D
[…]
A Standard diffusion loss 25 steps 24.8 0.891 0.215 22.1 0.870 0.267 36.7 0.994 0.011 30.1 0.988 0.024
B Standard diffusion loss 1 step 25.7 0.901 0.242 23.0 0.877 0.320 38.5 0.995 0.012 31.5 0.990 0.030
C $L_{latent}$ 1 step 25.6 0.901 0.238 22.9 0.875 0.318 38.3 0.995 0.012 31.4 0.990 0.030
D $L_{latent}+L_{\text{LPIPS}}+L_{\text{L1}}$ 1 step 26.2 0.915 0.179 22.8 0.886 0.244 38.2 0.995 0.009 30.9 0.989 0.024
Table 7: We analyse the impact of different training strategies and losses on the Stereo4D and Ego4D datasets. Metrics are reports over the full image, as well as over only the disoccluded regions.
Appendix C Implementation Details
C.1 Conditioning
We extend the first convolution of SVD that initially took 8-dimensional input (4 dim noise latent + 4 dim VAE encoded image) to the 13-dimensional input (4 dim noise latent + 4 dim VAE encoded left to view video + 4 dim VAE encoded warped view video + 1 dim disocclusion mask). To ensure smooth model initialization we copy weights 5-8 channels to 9-12 channels and divide both by factor 2 and initialize channels of dissolution mask with 0.
Model training:  We initialize our M2SVid model using the Stable Video Diffusion model [4]. We train the model for 300k iterations with a batch size of 16 and a learning rate of $2\times 10^{-6}$. Due to GPU memory limitations, during training, we sample batches with frame number and resolutions of $4\times 512\times 512$, $16\times 256\times 256$, and $25\times 192\times 192$, utilize gradient checkpointing[6] and train the model in float16 precision. We employ random resized cropping, using a resize scale in the range of [0.3, 1.0], for data augmentation. We preprocess the disocclusion masks using the closing morphological transformation to fill small holes inside the inpainting regions with a kernel size of 11 during training and testing.
Depth estimation with DepthCrafter:  During inference for depth prediction, we utilize the DepthCrafter [22] method, the state-of-the-art in diffusion-based video-depth estimation approach. This model can accommodate up to 110 frames and out temporally consistent depth. However, any depth model might be used at this step, for example, efficient 1-step depth diffusion models as in [16]. We predict depth in chunks of 110 frames with an overlap of 25 frames. Depth was predicted on frames resized to a resolution of 1024, followed by upscaling to the original size.
Appendix D More Details on Results
D.1 State-of-the-art Methods
SVG is a training-free method that utilizes a frozen diffusion model to inpaint regions within the mask while preserving the other regions. To achieve spatial-temporal consistency, the method employs inpainting of 8 uniformly sampled views between the given left view and the required right view at the same time, increasing the computational burden. We use 50 denoising steps as recommended. SVG [10] has been compared to state-of-the-art video inpainting [93, 38] and dynamic novel view synthesis [45, 39] methods, significantly outperforming them in temporal consistency and overall inpainting quality. This highlights the fundamental differences between stereo video inpainting, classical video inpainting, and novel view synthesis, revealing a substantial gap between these tasks. Due to this disparity, we exclude classical video inpainting and novel view synthesis methods from our comparisons.
StereoCrafter [68] is a concurrent method and the closest baseline to our approach. It fine-tunes a diffusion model, SVD, to perform inpainting given a warped view and a mask in a diffusion-based denoising setup, trained on a privately collected internet dataset. Since the method does not specify the number of denoising steps, we use 25 steps, as is commonly done [22].
Note that our method M2SVid was trained on the Stereo4D and Ego4d datasets, while SVG is a training-free method, and StereoCrafter is trained on a private dataset. However, this StereoCrafter private dataset also contains general internet videos and should have a similar data distribution. Comparing zero-shot performance on another dataset is challenging, as there are no publicly available stereo video datasets in the general domain. For this reason, we evaluate on the Stereo4D dataset.
D.2 User Study
In order to perform our human perception user study to rank the different methods, we selected a set of 21 videos. We tried to ensure a diverse collection of videos covering different types of subjects, video quality, and motion pattern. We avoided including without with very large motions since we found it very hard to compare the methods in such cases.
During the study, the participants were provided HTML pages containing the links to the generated predictions, as shown in Figure 7. Furthermore, the users were shown detailed instructions about the study at the top of the HTML page. The exact instructions are shown in Figure 8. In particular, the users were provided a brief background about the problem statement and the method. They were asked to ranking the quality of the different methods based on temporal consistency, image quality (i.e. sharpness), lack of artifacts, etc. In order to miminize noise, the users were allowed to assign same ranks to two or more methods in case they were indistinguishable.
We had 13 distinct participants in the study, out of which 9 were male and 4 were female. We divided the 21 videos into sets of 7 each. Each user was given the option of assigning rankings to the videos in one or two sets each. In total, we received rankings for 16 sets, resulting in 112 sets of ranks.
D.3 State-of-the-art Comparison
For quantitative evaluation, we use PSNR, MS-SSIM, and LPIPS [89] metrics, computed independently for each video and averaged over the full dataset. Additionally, we report these metrics separately for the disoccluded and non-disoccluded regions in Tab. 5. To achieve this, we mask all pixels outside the considered region with white pixels, compute the metrics at the video level, and then average them over the dataset.
Our method obtains the best PSNR and LPIPS scores when averaged over the full image. For the disoccluded regions, we see that single step StereoCrafter obtains the best PSNR score, while also having the worse LPIPs score. We believe this is because the method generates blurry results, which is often favored by PSNR, but strongly penalized by LPIPS. Outside the discocclusion region, our method obtains the best LPIPS and PSNR scores while SVG obtains the best MS-SSIM score.
D.4 Ablations
In Tabs. 6 and 7, we further report metrics for the disoccluded regions for our ablation studies. We find that incorporating the left view as an additional conditioning signal (Tab. 6) results in a slight improvement in inpainting performance across both datasets. Additionally, end-to-end training with latent space supervision (Tab. 7) achieves nearly the same inpainting performance as a model trained with the standard diffusion paradigm and evaluated with 1-step inference. However, end-to-end training with image-based loss leads to a substantial drop in LPIPS, particularly on the Ego4D dataset.
D.5 Run-time
Method Latency (s) $\downarrow$
without full attention on the disoccluded regions 2.0
with full attention on the disoccluded regions 2.1
Table 8: The latency of two variations of our model: with and without full attention on the disoccluded regions. Enabling full attention on the disoccluded regions results in only a slight increase in runtime. The run-times computed using an A100 GPU on a $512\times 512$ videos with 16 frames.
We compare the runtime of our model with and without full attention in the disoccluded regions in Tab. 8. Our results show that enabling full attention leads to only a slight increase in runtime.
Appendix E Inference on Arbitrary Length and Resolution
To achieve stereo video conversion for videos of arbitrary length, we fine-tune our M2SVid model for 20K iterations using the auto-regressive modeling approach proposed in [68]. Specifically, during training, we replace the first $n$ frames of $V_{r}^{warp}$ with ground truth frames from the right video $V_{r}$, where $n$ is randomly sampled between 00 and $N$. During inference on long videos, we use the last $m$ frames generated in the current round as input for the next round, ensuring seamless stitching between consecutive rounds. For inference with arbitrary resolution, following [68], we utilize tiled diffusion. Specifically, we divide high-resolution videos into overlapping blocks and perform stereo video conversion on each block independently. The overlapping regions are blended in the latent space before VAE decoding to ensure seamless transitions between blocks. We provide examples generated by our model on high-resolution long videos in the supplementary material zip file.